About 3 years ago I bought (very cheaply) a defunct HP 1725A oscilloscope. This was never intended to be my main oscilloscope, but I have great respect for well made vintage test gear, and you don't get much better than HP of yesteryear.
In its day this would have been pretty high end, and with a 275Mhz bandwidth and other features (e.g. 50 ohm input capability) it would still be of use to me.
Anyway I got it working of sorts, the main fault is that many of the contacts in the timebase switch wafers are broken. If I recall there are 5 (maybe 4?) such boards on the TB switch. I managed to kind of repair them with phosphor bronze strip - see photo - but switching TB is still unreliable. When going, it works pretty much perfectly, voltage and timebase very accurate.
Long shot - I'm wondering if anyone else has ever fixed one of these, especially the above mentoned switches, and has any tips or advice. As for actually getting spares, I think I've more chance of winning the lottery.
Many thanks
Charles